In any given month, Mark can produce 58 baseballs and Katie can produce 25 baseballs. Alternatively,...

  1. In any given month, Mark can produce 58 baseballs and Katie can produce 25 baseballs. Alternatively, Mark can produce 94 bats and Katie can produce 15 bats.

    What is Mark’s opportunity cost of producing 30 bats?

    Round your answer to two (2) decimal places
Answer #1
Baseballs Bats
Mark 58 94
Katie 25 15

Mark's opportunity cost of producing 1 bat = (58/94)= 0.617 baseballs

So, Mark's opportunity cost of producing 30 bats = (0.617)(30)= 18.51 baseballs.